OpenStudy (anonymous):

If I had a linear equation of -8x-8y=0 and my A points are 0,0 and if I used 1 for the value of y would my B points be 1, 16

OpenStudy (anonymous):

i dont understand ur ques

OpenStudy (anonymous):

your equation is -8x-8y=0 -8x = 8y -x=y so if y =1 x =-1

OpenStudy (anonymous):

Thank you